Hi
Joshua and his sister spent $96 in total. His sister spent 40% less than him. How much did he spend.
My solution was as follows
Joshua = 1.4x
Sister = x
2.4x = 96 x=40 therefore Joshua spent 56 dollars . The textbook says 60 dollars.
Who is correct. Thanks

Answer by ikleyn(52776)

.

Let x be the amount Joshua spent.


Then his sister spent 100% - 40% = 60% of x,  or  0.6x.


The equation for the total is


    x + 0.6x = 96

    1.6x     = 96

       x     = 96/1.6 = 60.


ANSWER.  Joshua spent 60 dollars.

Solved (correctly).

--------------

Your solution is  INCORRECT.   The textbook is  RIGHT.


    Why and where is your solution incorrect ?


    When you take x = what the sister spent, you INCORRECTLY assume/conclude that Joshua spent 1.4x.


    The condition says that  THE SISTER SPENT 40% LESS THAN JOHUA.


    It DOES NOT SAY that JOHUA SPENT 40% MORE THAN THE SISTER.


    These are two DIFFERENT statements: my interpretation IS CORRECT, while yours IS NOT.


    When you solve such problems, be VERY attentive and pay SPECIAL attention to identify WHAT IS THE BASE for comparison (!)


    From the context, in this problem the  JOSHUA' SPENT AMOUNT  is the base for the comparison - NOT the SISTER's amount (!)


Is everything clear to you in my solution/explanation ?

If  "YES",  then do not forget to post your  "THANKS"  for my teaching . . .